A wildebeest and a chicken participate in a race over a 1.80 km long course. The wildebeest travels at a speed of 20.0 m/s and the chicken can do 3.90 m/s. The wildebeest runs for 1.440 km and then stops to tease the slow-moving chicken, which eventually passes by. The wildebeest waits for a while after the chicken passes and then runs toward the finish line. Both animals cross the finish line at the exact same instant. Assume both animals, when moving, move steadily at their respective speeds. (a) How far (in m) is the chicken from the finish line when the wildebeest resumes the race? (b) For how long in time (in s) was the wildebeest stationary?
Answer #1

a)

D = total distance of the race = 1.80 km = 1800 m

v = speed at which chicken travels = 3.90 m/s

V = speed at which wildbeast travels = 20 m/s

d = distance traveled by wildbeast before stopping to tease = 1.440 km = 1440 m

tc = time taken by chicken to complete the race

time taken by chicken to complete the race is given as

tc = D/v = 1800/3.90 = 461.54 sec

ts = time taken by the wildbeast to travel the distance before stopping to tease

time taken by the wildbeast to travel the distance before stopping to tease is given as

ts = d/V = 1440/20 = 72 sec

Time taken by wildbeast to complete the race without stopping is given as

T = D/V = 1800/20 = 90 sec

time taken by the wildbeast to travel the remaining distance after stopping to tease is 90 - 72 = 18 sec.

Since both reach the finish line at the same time , hence chicken also travels for 18 sec to reach the finish line

t' = 18 sec

dchicken = distance of chicken from finishing line when the wildbeast starts again = v t'

dchicken = 3.90 x 18

dchicken = 70.2 m

b)\

tstationary = time for which the wildbeast was stationary = ?

tstationary = tc - T

tstationary = 461.54 - 90

tstationary = 371.54 sec